OpenStudy (javk):

The cubic polynomial x^3-2x^2-x-6 is denoted by f(x). Show that (x-3) is a factor of f(x). Factorize F(x). Hence find the number of real roots of the equation f(x) = 0. Justifying your answer. Hence write down the number of points of intersection of the graphs with equations y = x^2-2x-1 and y = 6/x justifying your answer.

OpenStudy (javk):

after factorization I got \[(x-3)(x ^{2}+x+2)\] there is only one real root. I am having trouble with the last part of the question where I find the number of points of intersection between the graphs. how should I proceed when the question says `hence`

OpenStudy (mww):

The second part, hence indicates your result should lead into the first question. Remember for two equations to intersect, you solve simultaneously and the solutions. y = x^2-2x-1 and y = 6/x x^2-2x-1 = 6/x This eqn is readily convertible into a cubic. See that the original equation is restored. Just note the solutions to x^2-2x-1 = 6/x will obviously not include x = 0, but otherwise are the same solutions as that of x^3-2x^2-x-6